Hi all,
I need some help with school study work. I need to solve this quadratic equation by using the square root property?
1: 4x^2 - 48= 0
2: I need to solve this quadratic equation by completing square?
x^2 + 24 = -10x

Anyone, please help me

Answer by fractalier(6550)

Okay, from 4x^2 - 48 = 0, we add 48 to both sides and then divide by four, giving us x^2 = 12. Take the root of both sides to give x = ± radical 12 = ± 2 radical 3.
The second problem, x^2 + 24 = -10x, requires us to move the -10x to the left side and the 24 to the right to get
x^2 + 10x = -24
Now we take half of the 10 and square it to get 25. That is what we add to both sides and get
x^2 + 10x + 25 = 1
(x + 5)^2 = 1
x + 5 = ±1
x = -5 ± 1 = -6 or -4